OpenStudy (anonymous):

the length of a rectangle is 3in longer than the width and that the perimeter of the rectangle is 78. Set up an equation involving only W, the width of the rectangle

OpenStudy (aravindg):

\[\huge \text{WELCOME TO OPENSTUDY!}\]

OpenStudy (anonymous):

thank you

OpenStudy (ajprincess):

Let the width of the rectangle be w and length be l. l=w+3 Perimeter=2(l+w) 78=2(w+3+w) Try simplfyng this and u will get the required equation. Can u do it @nburgher?

OpenStudy (anonymous):

the width would be 18 and length is 21

OpenStudy (ajprincess):

ya u r right:) bt u need to write only the equation

OpenStudy (anonymous):

ok. Using the same length from above, find a new perimeter if the new width is 4 more than one-third of the length

OpenStudy (ajprincess):

length is 21. what is 1/3 of it?

OpenStudy (ajprincess):

(1/3)*21=?

OpenStudy (anonymous):

7

OpenStudy (ajprincess):

new width is 4 more than 7. so what will be it?

OpenStudy (anonymous):

so the equation would be P=2(21)+2(7+4)?

OpenStudy (ajprincess):

ya:)

OpenStudy (anonymous):

thank you very much. I have been stuck on this for days

OpenStudy (ajprincess):

welcome:)